# Genealogies of records of stochastic processes with stationary increments as unimodular trees

### v1

@@ 1 Introduction
Concenring the last question on record representation of graphs, it is shown that under a mild assumption, certain classes of ordered directed trees can be represented as the component of 00 in the record graph of some stationary sequence. This includes the infinite trees associated to virtually all dynamics on ordered networks. These results are now stated more precisely after introducing some terminology from [5].
@@ 1.1.1 Foil classification theorem of unimodular networks
Components of either class $\mathcal{I}/\mathcal{I}$ or of class $\mathcal{I}/\mathcal{F}$ are (non-rooted) EFTs. The class $\mathcal{I}/\mathcal{I}$ corresponds to the component being infinite and all foils of the component being infinite. Further, every EFT of this class is characterised by the properties that it has one end and that all of its vertices have finitely many descendants. The class $\mathcal{I}/\mathcal{F}$ corresponds to the component being infinite and all foils of the component being finite. Every EFT of this class has a unique bi-infinite directed path, characterised by the set of vertices with infinitely many descendants. The remaining vertices have finitely many descendants.
@@ 2.2 Unimodularity and vertex-shifts
A covariant subset is a collection of subsets $\mathfrak{S}_{G}$ indexed by (locally-finite) connected networks $G$, where $\mathfrak{S}_{G}$ is a subset of $V(G)$, that satisfy the following properties; (1) covariance property: $\alpha(\mathfrak{S}_{G})=\mathfrak{S}_{\alpha(G)}$ for every network isomporphism $\alpha$, (2) measurability: the map $[G,o]\mapsto\mathbf{1}\{o\in\mathfrak{S}_{G}\}$ is measurable. A covariant partition is a collection of partitions $\Pi_{G}$ indexed by (locally-finite) connected networks $G$, where $\Pi_{G}$ is a partition of $V(G)$ and the collection satisfies the following two conditions, (1) every network isomorphism $\alpha$ satisfies $\alpha\circ\Pi_{G}=\Pi_{\alpha(G)}$, (2) the subset $\{[G,o,u]:u\in\Pi_{G}(o)\}$ is a measurable subset, where $\Pi_{G}(o)$ is the element that contains $o$.
@@ 2.3 Unimodular Eternal Galton-Watson Trees
Note the emphasis on the Family Tree being eternal in Theorem 11. For a similar characterisation result about finite Family Trees, see Proposition 50.
@@ 2.4 Typical re-rooting operation
$=\frac{1}{\mathbb{E}[\#V(\mathbf{T})]}\mathbb{E}\left[\sum_{v\in V(\mathbf{T})}h_{-}([\mathbf{T},v])\right]$
$=\mathbb{E}\left[h_{-}([\mathbf{T}^{\prime},\mathbf{o}^{\prime}])\right].$
@@ 4 Record vertex-shift on skip-free to the left random walks
This section is organised as follows. The first subsection describes the relation between the number of offsprings of a vertex in the record graph and the increment associated to it, the second subsection focuses on the properties of skip-free to the left random walks, which will be used to compute the distributions of the record graphs. Two new families of unimodular random rooted networks are introduced in the third subsection. Finally, the fourth subsection focuses on the aforementioned main objective.
@@ 4.3.1 Typically rooted Galton-Watson Tree (TGWT)
A construction for the modified subcritical Galton-Watson Tree $\mathfrak{T}$, a random tree that resembles $TGWT$, is given by Jonsson and Stefánsson [14] (for a simpler construction, see the subcritical case of [13, Section 5]). In their construction, a vertex can have infinite degree with positive probability and the tree is undirected. The special vertices of $\mathfrak{T}$, which belong to the spine, have the offspring distribution same as that of the ancestors of the root of $TGWT$, except for the last special vertex of $\mathfrak{T}$ which may have infinite children. Note that the undirected tree obtained by forgetting the directions of edges of $TGWT$ is not the same as $\mathfrak{T}$.
@@ 4.3.2 Bi-variate Eternal Kesten Tree
The unimodularised version of this tree is an instance of a unimodular tree belonging to the class $\mathcal{I}/\mathcal{F}$. Its constuction is given using the typically re-rooted joining operation (as in Def. 14). See Figure 6 for an illustration.
The next paragraph introduces the bi-variate Eternal Kesten Tree, denoted as $EKT(\alpha,\beta)$, with the offspring distributions $\alpha$ and $\beta$. This tree is unordered, unlabelled and parametrized by $\alpha$ and $\beta$. Although $EKT(\alpha,\beta)$ is not unimodular, a unimodularised version can be obtained from it using the typically re-rooted joining operation. To use this operation, it is necessary to have $m(\beta)<1$. However, to define $EKT(\alpha,\beta)$, it is sufficient to have $m(\alpha)<\infty$ and $m(\beta)\leq 1$.
@@ 4.4.1 Negative mean
The third objective is to give a characterisation of $TGWT$, which is detailed in Proposition 50.
@@
The next proposition gives a characterizing condition for a Typically rooted Galton-Watson Tree. The proof of this proposition depends on the following lemma which says that every unimodular tree is characterised by the descendant tree of its root.
@@
for any such event $A$ of the form given by Eq (25). Assume further that the events $A_{1},A_{2},\cdots,A_{k}$ depend only up to the $n$-th generation of $o$, which is the union of $D_{1}(o)$, $D_{2}(o)$, $\cdots$, $D_{n}(o)$. It suffices to show Eq. (26) for such events because of the local topology.
@@ 4.4.3 Positive mean
Recall the construction bi-variate Eternal Kesten Tree (see Subsection 4.3.2). Let $[\mathbf{T}_{0},\mathbf{o}_{0}]$ be a Family Tree whose distribution is given by the following: the offspring distribution of $\mathbf{o}_{0}$ is $\bar{\pi}$. The descendant trees of the children of $\mathbf{o}_{0}$ are independent Galton-Watson Trees with offspring distribution $\tilde{\pi}$, and they are independent of $d_{1}(\mathbf{o}_{0})$. Let $([\mathbf{T}_{i},\mathbf{o}_{i}])_{i\in\mathbb{Z}}$ be an i.i.d. sequence of Family Trees. Recall that the distribution of the Family Tree obtained from the typically re-rooted joining of $([\mathbf{T}_{i},\mathbf{o}_{i}])_{i\in\mathbb{Z}}$ is the unimodularised $EKT(\bar{\pi},\tilde{\pi})$. Since the mean $m(\tilde{\pi})<1$, the typically re-rooted joining operation is well-defined for this sequence. Recall the ECS order on the joining of $([\mathbf{T}_{i},\mathbf{o}_{i}])_{i\in\mathbb{Z}}$ (see Subsection 4.3.2).
@@
Figure 7: Illustration of the subtrees of ancestors of 00 on a trajectory conditioned on the event that $S_{n}\leq 0$ for all $n\leq 0$ when $\mathbb{E}[X_{0}]>0$. Each box represents the subtrees $\mathbf{T}_{1},\mathbf{T}_{2},\mathbf{T}_{3},\cdots$ (see Step 1 of Theorem 53).
@@ 5 Record representable EFTs
The sequence $y=(y_{n})_{n\in\mathbb{Z}}$ takes values in $\{-1,0,1,2,\cdots\}^{\mathbb{Z}}$. It is shown that $\Psi_{R}\circ\Phi_{R}=I$ on $\widetilde{\mathcal{T}}_{*}$, where $\Psi_{R}([\mathbb{Z},0,y])=[\mathbb{Z}^{R}_{y}(0),0]$ is the component of 00 in the record graph of the network $(\mathbb{Z},y)$. The proof relies on the property that every elder sibling of any vertex (if they exist) of a Family Tree with a unique succession line has finitely many descendants.
@@
It is now shown that the sequence $Y$ is stationary. Let $(u_{n})_{n\in\mathbb{Z}}=U((\mathbf{T},\mathbf{o}))$ be the succession line with $u_{0}=\mathbf{o}$. Since $a$ and $b$ are well-defined on a unimodular EFT (by Lemma 56), it implies that the map $u_{n}\mapsto u_{n+i},\forall n\in\mathbb{Z}$ is bijective for any $i\in\mathbb{Z}$. Therefore, the map $[\mathbf{T},\mathbf{o}]\mapsto[\mathbf{T},u_{i}]$ is measure-preserving, for any $i\in\mathbb{Z}$, by Point-stationarity Theorem 5. Since, $[\mathbb{Z},0,T_{i}Y]=\Phi_{R}([\mathbf{T},u_{i}])$, the map $(Y_{n})_{n\in\mathbb{Z}}\mapsto(Y_{n+i})_{n\in\mathbb{Z}}$ is measure-preserving for any $i\in\mathbb{Z}$,. Thus, $Y$ is stationary. ∎
@@ 5.1 Examples of record representable vertex-shifts
for all $i\in\mathbb{Z}$. Let $X=(X_{n})_{n\in\mathbb{Z}}$ be an i.i.d. sequence of random variables taking values in $\mathbb{Z}$ such that their common mean exists. Let $\Psi_{SR}$ be the map that takes any network of the form $[\mathbb{Z},0,x]$ and maps to the component of 00 in the $SR$-graph of $(\mathbb{Z},x)$. One can show that the component of 00 in the $SR$-graph is of class $\mathcal{I}/\mathcal{I}$ if $\mathbb{E}[X_{0}]=0$ and $\mathbb{E}[X_{0}^{2}]>0$, because $\limsup_{n\to\infty}S_{-n}=\infty$; which implies that the descendant tree of 00 is a.s. finite, and because 00 has infintely many ancestors; which follows from the Chung-Fuchs recurrence theorem. By Theorem 54, there exists a stationary sequence $Y=(Y_{n})_{n\in\mathbb{Z}}$ of random variables taking values in $\{-1,0,1,2,\cdots\}$ such that $[\mathbb{Z}^{SR}_{X}(0),0]\overset{\mathcal{D}}{=}[\mathbb{Z}_{Y}^{R}(0),0]$.
Figure 9: An Illustration of Example 62. The top left figure shows the trajectory of a realisation of the i.i.d. sequence $X=(X_{n})_{n\in\mathbb{Z}}$, where $X_{n}\in\mathbb{Z},\forall n$. The $SR$-graph (strict record graph) is drawn in red on the trajectory of the top left figure. The component of 00 in the $SR$-graph is drawn in the bottom figure. The top right figure shows the trajectory of the sequence $Y=(Y_{n})_{n\in\mathbb{Z}}$ which is obtained by applying the backward map $\Phi_{R}$ to the rooted tree of the bottom figure. The component of 00 in the record graph of the trajectory in the top right figure is the same as the component of 00 in the $SR$-graph.
@@
Figure 10: An Illustration of Example 63. The top figure shows the trajectory of a realisation of the i.i.d. sequence $X=(X_{n})_{n\in\mathbb{Z}}$, where $X_{n}\in\mathbb{Z},\forall n$. The $C$-graph ($C$ for climbing point vertex-shift) is drawn in red on the trajectory of the top figure. The vertices on the bi-infinite path are highlighted. The component of 00 in the $C$-graph is drawn in the middle figure. The bottom figure shows the trajectory of the sequence $Y=(Y_{n})_{n\in\mathbb{Z}}$ which is obtained by applying the backward map $\Phi_{R}$ to the rooted tree shown in the middle figure. The component of 00 in the record graph of the trajectory in the bottom figure is the same as the component of 00 in the $C$-graph.

### v2

@@ 1 Introduction
Concerning the last question on record representation of graphs, it is shown that under a mild assumption, certain classes of ordered directed trees can be represented as the component of 00 in the record graph of some stationary sequence. This includes the infinite trees associated to virtually all dynamics on ordered networks, thereby underscoring the importance of delving into the record graphs of stationary sequences.
These results are now stated more precisely after introducing some terminology from [5].
@@ 1.1.1 Foil classification theorem of unimodular networks
Components of either class $\mathcal{I}/\mathcal{I}$ or of class $\mathcal{I}/\mathcal{F}$ are (non-rooted) EFTs. The class $\mathcal{I}/\mathcal{I}$ corresponds to the component being infinite and all foils of the component being infinite. Further, every EFT of this class is characterized by the properties that it has one end and that all of its vertices have finitely many descendants. The class $\mathcal{I}/\mathcal{F}$ corresponds to the component being infinite and all foils of the component being finite. Every EFT of this class has a unique bi-infinite directed path, characterized by the set of vertices with infinitely many descendants. The remaining vertices have finitely many descendants.
@@ 2.2 Unimodularity and vertex-shifts
A covariant subset is a collection of subsets $\mathfrak{S}_{G}$ indexed by (locally-finite) connected networks $G$, where $\mathfrak{S}_{G}$ is a subset of $V(G)$, that satisfy the following properties; (1) covariance property: $\alpha(\mathfrak{S}_{G})=\mathfrak{S}_{\alpha(G)}$ for every network isomorphism $\alpha$, (2) measurability: the map $[G,o]\mapsto\mathbf{1}\{o\in\mathfrak{S}_{G}\}$ is measurable. A covariant partition is a collection of partitions $\Pi_{G}$ indexed by (locally-finite) connected networks $G$, where $\Pi_{G}$ is a partition of $V(G)$ and the collection satisfies the following two conditions, (1) every network isomorphism $\alpha$ satisfies $\alpha\circ\Pi_{G}=\Pi_{\alpha(G)}$, (2) the subset $\{[G,o,u]:u\in\Pi_{G}(o)\}$ is a measurable subset, where $\Pi_{G}(o)$ is the element that contains $o$.
@@ 2.3 Unimodular Eternal Galton-Watson Trees
Note the emphasis on the Family Tree being eternal in Theorem 11. For a similar characterization result about finite Family Trees, see Proposition 50.
@@ 2.4 Typical re-rooting operation
$=\frac{1}{\mathbb{E}[\#V(\mathbf{T})]}\mathbb{E}\left[\sum_{v\in V(\mathbf{T})}h_{-}([\mathbf{T},v])\right]=\mathbb{E}\left[h_{-}([\mathbf{T}^{\prime},\mathbf{o}^{\prime}])\right].$
@@ 4 Record vertex-shift on skip-free to the left random walks
This section is organized as follows. The first subsection describes the relation between the number of offsprings of a vertex in the record graph and the increment associated to it, the second subsection focuses on the properties of skip-free to the left random walks, which will be used to compute the distributions of the record graphs. Two new families of unimodular random rooted networks are introduced in the third subsection. Finally, the fourth subsection focuses on the aforementioned main objective.
@@ 4.3.1 Typically rooted Galton-Watson Tree (TGWT)
A construction for the modified subcritical Galton-Watson Tree $\mathfrak{T}$, a random tree that resembles $TGWT$, is given by Jonsson and Stefánsson [15] (for a simpler construction, see the subcritical case of [14, Section 5]). In their construction, a vertex can have infinite degree with positive probability and the tree is undirected. The special vertices of $\mathfrak{T}$, which belong to the spine, have the offspring distribution same as that of the ancestors of the root of $TGWT$, except for the last special vertex of $\mathfrak{T}$ which may have infinite children. Note that the undirected tree obtained by forgetting the directions of edges of $TGWT$ is not the same as $\mathfrak{T}$.
@@ 4.3.2 Bi-variate Eternal Kesten Tree
The unimodularised version of this tree is an instance of a unimodular tree belonging to the class $\mathcal{I}/\mathcal{F}$. Its construction is given using the typically re-rooted joining operation (as in Def. 14). See Figure 6 for an illustration.
The next paragraph introduces the bi-variate Eternal Kesten Tree, denoted as $EKT(\alpha,\beta)$, with the offspring distributions $\alpha$ and $\beta$. This tree is unordered, unlabeled and parametrized by $\alpha$ and $\beta$. Although $EKT(\alpha,\beta)$ is not unimodular, a unimodularised version can be obtained from it using the typically re-rooted joining operation. To use this operation, it is necessary to have $m(\beta)<1$. However, to define $EKT(\alpha,\beta)$, it is sufficient to have $m(\alpha)<\infty$ and $m(\beta)\leq 1$.
@@ 4.4.1 Negative mean
The third objective is to give a characterization of $TGWT$, which is detailed in Proposition 50.
@@
The next proposition gives a characterizing condition for a Typically rooted Galton-Watson Tree. The proof of this proposition depends on the following lemma which says that every unimodular tree is characterized by the descendant tree of its root.
@@
for any such event $A$ of the form given by Eq (25). Assume further that the events $A_{1},A_{2},\cdots,A_{k}$ depend only on the descendant tree up to the $n$-th generation of $o$, which is the union of $D_{1}(o)$, $D_{2}(o)$, $\cdots$, $D_{n}(o)$. It suffices to show Eq. (26) for such events because of the local topology.
@@ 4.4.3 Positive mean
Recall the construction bi-variate Eternal Kesten Tree (see Subsection 4.3.2). Let $[\mathbf{T}_{0},\mathbf{o}_{0}]$ be a Family Tree whose distribution is given by the following: the offspring distribution of $\mathbf{o}_{0}$ is $\bar{\pi}$. The descendant trees of the children of $\mathbf{o}_{0}$ are independent Galton-Watson Trees with offspring distribution $\tilde{\pi}$, and they are independent of $d_{1}(\mathbf{o}_{0})$. Let $([\mathbf{T}_{i},\mathbf{o}_{i}])_{i\in\mathbb{Z}}$ be an i.i.d. sequence of Family Trees. Recall that the distribution of the Family Tree obtained from the typically re-rooted joining of $([\mathbf{T}_{i},\mathbf{o}_{i}])_{i\in\mathbb{Z}}$ is the unimodularised $EKT(\bar{\pi},\tilde{\pi})$. Since the mean $m(\tilde{\pi})<1$ (by Lemma 42), the typically re-rooted joining operation is well-defined for this sequence. Recall the ECS order on the joining of $([\mathbf{T}_{i},\mathbf{o}_{i}])_{i\in\mathbb{Z}}$ (see Subsection 4.3.2).
@@
Figure 7: Illustration of the subtrees of ancestors of 00 on a trajectory conditioned on the event that $S_{n}\leq 0$ for all $n\leq 0$ when $\mathbb{E}[X_{0}]>0$. Trajectories in the boxes correspond to the subtrees $\mathbf{T}_{1},\mathbf{T}_{2},\mathbf{T}_{3},\cdots$ (see Step 1 of Theorem 53).
@@
Recall the construction of unimodularised bi-variate $EKT(\alpha,\beta)$. Note that the offspring distributions $\alpha,\beta$ need to have means $m(\alpha)<\infty$ and $m(\beta)<1$ to define the typically re-rooted joining of i.i.d. bi-variate $GW(\alpha,\beta)$ Family Trees. It is now shown that $m(\bar{\pi})<\infty$ (since $\alpha=\bar{\pi}$ here). This follows from the bound $X_{\tau-1}-S_{\tau}\leq\tau$, which is a consequence of the skip-free to the left property. Since $\mathbb{E}[X_{0}]>0$, the classical result on stopping time [13] implies $\mathbb{E}[\tau]<\infty$. Thus, $m(\bar{\pi})=\mathbb{E}[X_{\tau-1}-S_{\tau}]\leq\mathbb{E}[\tau]<\infty$. The fact that $m(\tilde{\pi})<1$ follows from Lemma 42.
@@ 5 Record representable EFTs
Example 59 (Unimodular EFT with two succession lines).
Consider the unimodularised bi-variate Eternal Kesten Tree with offspring distributions $\alpha,\beta$, as described in Subsection 4.3.2, but with a uniform order imposed on the vertices of the bi-infinite path instead of the ECS order. The construction of this tree is as follows: start with the typically re-rooted joining $[\mathbf{T},\mathbf{o}]$ of i.i.d. sequence of bi-variate ordered $GW(\alpha,\beta)$, where the offspring distributions $\alpha$ and $\beta$ have means $m(\alpha)<\infty$ and $m(\beta)<1$. Let $(\mathbf{o}^{\prime}_{n})_{n\in\mathbb{Z}}$ denote the vertices of the bi-infinite path of $[\mathbf{T}^{\prime},\mathbf{o}^{\prime}]$, where $(\mathbf{o}^{\prime}_{n})$ is the child of $(\mathbf{o}^{\prime}_{n+1})$, for all $n\in\mathbb{Z}$, and $\mathbf{o}_{0}^{\prime}$ is the ancestor of the root $\mathbf{o}^{\prime}$. For each $n\in\mathbb{Z}$, assign a uniform order to the children of $\mathbf{o}_{n}^{\prime}$. In particular, given that $\mathbf{o}_{n+1}^{\prime}$ has $k$ children, the order of $\mathbf{o}_{n}^{\prime}$ is uniformly distributed on ${1,2,\ldots,k}$. The resulting unimodularised ordered bi-variate $EKT(\alpha,\beta)$ $[\mathbf{T},\mathbf{o}]$ has infinitely many vertices in the set $W$ (Eq. 35) with positive probability. According to Lemma 57, $[\mathbf{T},\mathbf{o}]$ has two succession lines with positive probability.
@@
The sequence $y=(y_{n})_{n\in\mathbb{Z}}$ takes values in $\{-1,0,1,2,\cdots\}^{\mathbb{Z}}$. It is shown that $\Psi_{R}\circ\Phi_{R}=I$ on $\widetilde{\mathcal{T}}_{*}$, where $\Psi_{R}([\mathbb{Z},0,y])=[\mathbb{Z}^{R}_{y}(0),0]$ is the component of 00 in the record graph of the network $(\mathbb{Z},y)$. The proof relies on the property that every elder sibling of any vertex (if they exist), of a Family Tree with a unique succession line, has finitely many descendants.
@@
It is now shown that the sequence $Y$ is stationary. Let $(u_{n})_{n\in\mathbb{Z}}=U((\mathbf{T},\mathbf{o}))$ be the succession line with $u_{0}=\mathbf{o}$. Since $a$ and $b$ are well-defined on a unimodular EFT (by Lemma 56), it implies that the map $u_{n}\mapsto u_{n+i},\forall n\in\mathbb{Z}$ is bijective for any $i\in\mathbb{Z}$. Therefore, the map $[\mathbf{T},\mathbf{o}]\mapsto[\mathbf{T},u_{i}]$ is measure-preserving, for any $i\in\mathbb{Z}$, by Point-stationarity Theorem 5. Since, $[\mathbb{Z},0,T_{i}Y]=\Phi_{R}([\mathbf{T},u_{i}])$, the map $(Y_{n})_{n\in\mathbb{Z}}\mapsto(Y_{n+i})_{n\in\mathbb{Z}}$ is measure-preserving for any $i\in\mathbb{Z}$. Thus, $Y$ is stationary. ∎
@@ 5.1 Examples of record representable vertex-shifts
for all $i\in\mathbb{Z}$. Let $X=(X_{n})_{n\in\mathbb{Z}}$ be an i.i.d. sequence of random variables taking values in $\mathbb{Z}$ such that their common mean exists. Let $\Psi_{SR}$ be the map that takes any network of the form $[\mathbb{Z},0,x]$ and maps to the component of 00 in the $SR$-graph of $(\mathbb{Z},x)$. One can show that the component of 00 in the $SR$-graph is of class $\mathcal{I}/\mathcal{I}$ if $\mathbb{E}[X_{0}]=0$ and $\mathbb{E}[X_{0}^{2}]>0$, because $\limsup_{n\to\infty}S_{-n}=\infty$; which implies that the descendant tree of 00 is a.s. finite, and because 00 has infinitely many ancestors; which follows from the Chung-Fuchs recurrence theorem. By Theorem 54, there exists a stationary sequence $Y=(Y_{n})_{n\in\mathbb{Z}}$ of random variables taking values in $\{-1,0,1,2,\cdots\}$ such that $[\mathbb{Z}^{SR}_{X}(0),0]\overset{\mathcal{D}}{=}[\mathbb{Z}_{Y}^{R}(0),0]$.
Figure 9: An Illustration of Example 63. The top left figure shows the trajectory of a realization of the i.i.d. sequence $X=(X_{n})_{n\in\mathbb{Z}}$, where $X_{n}\in\mathbb{Z},\forall n$. The $SR$-graph (strict record graph) is drawn in red on the trajectory of the top left figure. The component of 00 in the $SR$-graph is drawn in the bottom figure. The top right figure shows the trajectory of the sequence $Y=(Y_{n})_{n\in\mathbb{Z}}$ which is obtained by applying the backward map $\Phi_{R}$ to the rooted tree of the bottom figure. The component of 00 in the record graph of the trajectory in the top right figure is the same as the component of 00 in the $SR$-graph.
@@
Figure 10: An Illustration of Example 64. The top figure shows the trajectory of a realization of the i.i.d. sequence $X=(X_{n})_{n\in\mathbb{Z}}$, where $X_{n}\in\mathbb{Z},\forall n$. The $C$-graph ($C$ for climbing point vertex-shift) is drawn in red on the trajectory of the top figure. The vertices on the bi-infinite path are highlighted. The component of 00 in the $C$-graph is drawn in the middle figure. The bottom figure shows the trajectory of the sequence $Y=(Y_{n})_{n\in\mathbb{Z}}$ which is obtained by applying the backward map $\Phi_{R}$ to the rooted tree shown in the middle figure. The component of 00 in the record graph of the trajectory in the bottom figure is the same as the component of 00 in the $C$-graph.